OpenStudy (gabylovesyou):

Write an equation of a line perpendicular to y = 1 over 4x + 4 in slope-intercept form that passes through the point (−2, 6). y= 1 over 4x − 4 y= 1 over 4x + 6 y= −4x − 2 y= −4x + 1 over 4

OpenStudy (gabylovesyou):

@Compassionate @jim_thompson5910

OpenStudy (dan815):

y=1/4 x + 4 is a line with slope 1/4 to be perpendicular you have to take negative reciprocal of the slope so -4/1 = -4 y=-4x+b, now we need to find the right intercept, so that it goes through the point -2,6

OpenStudy (dan815):

6=-4(-2)+b b=6-8=-2 so y=-4x-2

OpenStudy (gabylovesyou):

C... Thanks so much! Your explanation is great!